La bibliographie de la partie ”Méthode de Capon”

A. Coulais

April 4, 2003

References

Burg72 [1] J. Burg, The relationship between maximum entropy spectra and maximum likelihood spectra, Geophysics, 37(2):375– 376, Avril 1972, .

Cap69 [2] J. Capon, High-resolution frequency-wavenumber spectrum analysis, Proc. IEEE, 57(8):1408–1418, Août 1969, .

Cap67 [3] J. Capon, R. Greenfield, and R. Kolker, Multidimensional Maximum–Likelihood processing of a large aperture seismic array, Proc. IEEE, 55(2):192–211, Février 1967, .

EBMP76 [4] I. El-Behery and R. MacPhie, Radio source parameters estimation by Maximum–Likelihood processing of variable baseline correlation interferometer data, IEEE Trans. on antennas and propagation, 24(2):163–173, Mars 1976, .

EBMP77 [5] I. El-Behery and R. MacPhie, Maximum–Likelihood estimation of source parameters from time–sampled outputs of a linear array, Journal of the Acoustical Society of America, 62(1):125–134, Juillet 1977, .

EBMP78 [6] I. El-Behery and R. MacPhie, Maximum likelihood estimation of the number, directions and strengths of point radio sources from variable baseline interferometer data, IEEE Trans. on antennas and propagation, 26(2):234–238, Mars 1978, .

EBMP80 [7] I. El-Behery and R. MacPhie, The resolution limit of a variable baseline synthetic aperture antenna, IEEE Trans. on antennas and propagation, 28(2):234–238, Mars 1980, .

MAX87_Escu [8] B. Escudié, Application de la représentation Temps-Fréquence à l’imagerie acoustique spatio–fréquentielle, in Méthodes et techniques de traitement du signal et applications aux mesures physiques, J. Max and collaborateurs, eds., vol. 2, Masson, Paris, 1987, ch. XXVI, pp. 332–360.

HAYKIN85_Hay [9] S. Haykin, Radar array processing for angle of arrival estimation, in Array signal processing, S. Haykin, ed., Prentice-Hall, 1985, livre 4, pp. 194–292.

HAYKIN85_Jus [10] J. Justice, Array processing in exploration seismology, in Array signal processing, S. Haykin, ed., Prentice-Hall, 1985, livre 2, pp. 6–114.

KaMa81 [11] S. Kay and S. Marple, Spectrum analysis : a modern perspective, Proc. IEEE, 69(11):1380–1419, Novembre 1981, .

MAX87_Lac [12] J.-L. Lacoume, Nouvelles méthodes d’analyse spectrale, in Méthodes et techniques de traitement du signal et applications aux mesures physiques, J. Max and collaborateurs, eds., vol. 2, Masson, Paris, 1987, ch. XXI, pp. 131– 164.

Lagu86 [13] M. Lagunas, M. Santamaria, A. Gasull, and A. Moreno, Maximum likelihood filters in spectral estimation problems, Signal Processing, 10(0):19–34, mois inconnu 1986, .

Papa75 [14] G. Papadopoulos, A statistical technique for processing radio interferometer data, IEEE Trans. on antennas and propagation, 23(1):45–53, Janvier 1975, .

Park77 [15] R. Parker, Understanding inverse theory, Annual Revue of Earth and Planets Sciences, 5(0):35–64, mois inconnu 1977, .

PiLe91 [16] J.-L. Pinçon and f. Lefeuvre, Local characterization of homogeneous turbulence in a space plasma from simultaneous measurements of field components at several points in space, Journal of Geophysical Research, 96(A2):1789–1802, Février 1991, .